OpenStudy (fredoneedshelp):

Nine more than the product of a number and 2 is equal to 4

OpenStudy (jdcap):

\[x\cdot2+9=4\\ x=-\frac{ 5 }{ 2 }\]

OpenStudy (mysticfury):

-2.5 is your answer 2 x -2.5 = -5 -5 + 9 = 4

OpenStudy (fredoneedshelp):

Its just translating into a multi-step equation

OpenStudy (ellalaynemaguire):

9 + (2x) = 4

OpenStudy (qwertty123):

@JDcap @ellalaynemaguire @MysticFury Please try to avoid giving such direct answers, it is against OpenStudy's code of conduct http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. So, please keep that in mind from now on :)

OpenStudy (jdcap):

Ok, I haven't read that. Excuse me and thank you.

OpenStudy (qwertty123):

You Welcome ^.^